If the string is 9.5 m long, has a mass of 40 g and is pulled taut with a tension of 8.0 N , how much time does it take for a wave to travel from one end of the string to the other?
Answer #1

velocity = Sqrt (F/(m/L))

velocity = d/t

d/t = sqrt(F/(m/L))

t = d / (sqrt(F/(m/L))

time = 9.5 m / sqrt (8.0 N / (0.040 kg/9.5 m))

time = 2.07 s

Please Rate the answer !!!